Given that y = x is a particular solution of the differential

equation y''-x2y'

+ xy = 0. Find its general solution.



1
Expert's answer
2021-07-28T13:43:45-0400

We make the substitution "y=y_1u=xu"

Then

"y'=u+xu'"

"y''=u'+u'+xu''=2u'+xu''"

Substitute


"2u'+xu''-x^2u-x^3u'+x^2u=0"

"xu''+(2-x^3)u'=0"

Let "p=u'." Then "u''=p'."


"xp'+(2-x^3)p=0"

"\dfrac{dp}{p}=\dfrac{x^3-2}{x}dx"

Integrate both sides


"\int\dfrac{dp}{p}=\int\dfrac{x^3-2}{x}dx"

"\ln|p|=\dfrac{x^3}{3}-2\ln|x|+\ln C"


"p=\dfrac{c_1e^{x^3/3}}{x^2}"

"u=\int \dfrac{c_1e^{x^3/3}}{x^2}dx"

"y=x\int \dfrac{c_1e^{x^3/3}}{x^2}dx"